"i" companies operate in the frozen fish market, each of which has the following supply function Qis = 10+ P (i = 1,2,3 .., 100),
and there are 200 consumers each of whom has the following individual demand function Qid = 20-P (i = 1,2,3 ..., 200). Build supply and demand
From the market. Determine the equilibrium price and quantity. What happens if the price drops to half the equilibrium price? What happens if the price rises to half the
equilibrium price?
Answer #1

Supply of individual companies Qi - 10+P supply of 100 such companies (Market Supply Q = 100Q = 100 (roup) 1000 + 100 p dema1000+100P = 4000 - 200p Q* = 1000+ 100pt E 1000 + 100 x 10 a 1000 + 1ooo * - 2000 If price doops to half the equemprice new PQ - 3000 Since Q5<QP, there will be a shortage of frozen fish If price rises to half the equby price (Question is wrong. Pric